A solution contains 1.60×10-2 M zinc acetate and 1.60×10-2 M calcium nitrate. Solid potassium carbonate is added slowly to this mixture. What is the concentration of zinc ion when calcium ion begins to precipitate? Solubility product constant data is found in the Chemistry References.

[Zn2+] = M
Answer #1

When calcium ion begins to precipitate as calcium carbonate

Ca2+1 = Ca(NO3)2) = 1.60 x 10-2 M

The solubility product constant for calcium carbonate is Ksp = 3.3 x 10-4

The solubility product constant expression for calcium carbonate is Ksp = [Ca+] x [Co31

3.3 x 10-9 = 1.60 x 10-2 M x [Co-

3.3 x 10-9 CO3 1 = 1.60 x 10-2

3.3 x 10-9 CO3 1 = 1.60 x 10-2

CO-1= 2.06 x 10- M

When calcium ion begins to precipitate, the carbonate ion concentration is 2.06 x 10-7 M . Use this information and solubility product constant of zinc carbonate to calculate zinc ion concentration.

The solubility product constant for zinc carbonate is Ksp = 1.0 x 10-10

The solubility product constant expression for zinc carbonate is Ksp = [Zn2+] x [CO31

1.0 x 10-10 = Zn2+1 X 2.06 x 10-7 M

1.0 x 10-10 Zn2+] = 2.06 x 10-7

[Zn^{2+} ] =4.85 \times 10^{-4} \ M